Rover spends most of his time in a yard fenced by 54 meters of fencing. One side is 3 meters less than twice the length of the other. What are the dimensions of rovers yard?
Answer by Alan3354(69443)

Rover spends most of his time in a yard fenced by 54 meters of fencing. One side is 3 meters less than twice the length of the other. What are the dimensions of rovers yard?
=========
P = 2L + 2W = 54
L + W = 27
L = 2W-3
------
2W-3 + W = 27
etc
-------------------
PS If you don't like a dog in the house with you, don't get a dog.
Dogs are social animals. Getting just 1 dog and putting it outside is animal cruelty.